# CoCoA: A Generalized Approach to Uncertainty Quantification by Integrating Confidence and Consistency of LLM Outputs

### v2

@@ B.3 Different Ways of Combining Confidence and Consistency
$\mathrm{U}_{FullSampleCoCoA}=u_{*}^{\text{info}}\cdot U_{Deg}.$ (12)
@@
XSum WMT14FrEn WMT19DeEn CoQa Trivia MMLU GSM8k
@@
Table 8: Comparison of PRRs of CoCoA-family methods with alternative formulations with greedy sample taken in isolation.
@@
XSum WMT14FrEn WMT19DeEn CoQa Trivia MMLU GSM8k
@@
Table 9: Comparison of PRRs of CoCoA-family methods with alternative formulations of with best sample taken in isolation.
Appendix C Detailed Description of Uncertainty Quantification Methods
@@
C.1 Information-Based Methods
@@
Maximum Sequence Probability (MSP) is one of the simplest and most direct methods for estimating uncertainty. It measures the probability of the most likely output sequence given a specific input. Thus, uncertainty is quantified by calculating the probability of the sequence with the highest likelihood, under the assumption that the model is most confident in this output. It is defined as:
$U_{MSP}(\mathbf{y}\mid\mathbf{x},\bm{\theta})=-\log P(\mathbf{y}\mid\mathbf{x}).$ (13)
Perplexity (PPL) is another widely used metric for estimating uncertainty in language models (Fomicheva et al., 2020). It measures the model’s confidence by evaluating the average likelihood of generating the sequence tokens:
$U_{\mathrm{PPL}}(\mathbf{y},\mathbf{x})=-\frac{1}{L}\log P(\mathbf{y}\mid\mathbf{x}).$ (14)
Mean Token Entropy takes a broader view of uncertainty by considering the token-level predictions across the entire sequence (Fomicheva et al., 2020). Instead of evaluating the model’s confidence in a single output or individual token predictions, Mean Token Entropy calculates the average entropy of the token probability distributions for each token in the sequence:
$U_{\mathcal{H}_{T}}(\mathbf{y},\mathbf{x})=\frac{1}{L}\sum_{l=1}^{L}\mathcal{H}(y_{l}\mid\mathbf{y}_{<l},\mathbf{x}),$ (15)
@@
The TokenSAR method, introduced in (Duan et al., 2024), generalizes length-normalized log probability by computing a weighted average of the negative log probabilities of generated tokens, where weights are based on token relevance to the overall text. Using a similarity function $g(\cdot,\cdot)$ and token relevance function $R_{T}(y_{k},\mathbf{y},\mathbf{x})=1-g(\mathbf{x}\cup\mathbf{y},\mathbf{x}\cup\mathbf{y}\setminus y_{k})$, the uncertainty estimate is calculated as:
$U_{\mathrm{TokenSAR}}(\mathbf{y},\mathbf{x})=-\sum_{l=1}^{L}\tilde{\mathrm{R}}_{T}(y_{l},\mathbf{y},\mathbf{x})\log P(y_{l}\mid\mathbf{y}_{<l},\mathbf{x}),$ (16)
@@
$\tilde{\mathrm{R}}_{T}(y_{k},\mathbf{y},\mathbf{x})=\frac{\mathrm{R}_{T}(y_{k},\mathbf{y},\mathbf{x})}{\sum\nolimits_{l=1}^{L}\mathrm{R}_{T}(y_{l},\mathbf{y},\mathbf{x})}.$ (17)
@@
C.2 Consistency-Based Methods
@@
Consistency-based methods assess the uncertainty of a language model by evaluating the semantic consistency of its predictions across multiple outputs for the same prompt. The core idea is that semantically similar outputs indicate higher confidence, while diverse or conflicting outputs suggest greater uncertainty. Since language models can express the same meaning in different surface forms, these methods construct a semantic similarity matrix $G=(g_{ij})$, where each entry represents the degree of similarity between pairs of responses. By clustering responses into groups with equivalent meanings, these methods provide a semantic measure of the model’s consistency.
@@
Lin et al. (2024) offers two similarity measures to evaluate the similarity of sequences. The first is the Jaccard similarity, which treats sequences as sets of words and calculates the proportion of shared words to the total number of unique words in both sequences: $g(\mathbf{y},\mathbf{y}^{\prime})=|\mathbf{y}\cap\mathbf{y}^{\prime}|/|\mathbf{y}\cup\mathbf{y}^{\prime}|$.
@@
Natural Language Inference (NLI) provides another method for computing similarity between sequences. We use the DeBERTa-large NLI model (He et al., 2021), following Kuhn et al. (2023). For each pair of sequences, an NLI model predicts two probabilities: ${p}_{\mathrm{entail}}(\mathbf{y},\mathbf{y}^{\prime})$, indicating entailment, and ${p}_{\mathrm{contra}}(\mathbf{y},\mathbf{y}^{\prime})$, indicating contradiction. Similarity is then defined as either $g_{\mathrm{entail}}(\mathbf{y},\mathbf{y}^{\prime})={p}_{\mathrm{entail}}(\mathbf{y},\mathbf{y}^{\prime})$ or $g_{\mathrm{contra}}(\mathbf{y},\mathbf{y}^{\prime})=1-{p}_{\mathrm{contra}}(\mathbf{y},\mathbf{y}^{\prime})$.
@@
Among the simplest consistency-based approaches are the Number of Semantic Sets and the Sum of Eigenvalues of the Graph Laplacian (Lin et al., 2024). Number of Semantic Sets estimates how many distinct “meanings” the model produces by clustering its outputs with an NLI model. The number of semantic sets is initially equal to the total number of generated answers, $M$. Two sentences are grouped into the same cluster if the following conditions are satisfied: ${p}_{\text{entail}}(\mathbf{y}^{i},\mathbf{y}^{j})>{p}_{\text{contra}}(\mathbf{y}^{i},\mathbf{y}^{j})\quad\text{and}\quad{p}_{\text{entail}}(\mathbf{y}^{j},\mathbf{y}^{i})>{p}_{\text{contra}}(\mathbf{y}^{j},\mathbf{y}^{i})$. This computation is performed for all pairs of answers, and the final number of distinct clusters is denoted by $U_{\text{NumSemSets}}$.
@@
Sum of Eigenvalues of the Graph Laplacian examines global diversity: it constructs a similarity matrix among the sampled outputs and computes a continuous uncertainty score from the eigenvalues of the Laplacian of that similarity graph. Lin et al. (2024) proposes computing an averaged similarity matrix as $g_{ij}=\bigl{(}g\bigl{(}\mathbf{y}^{(i)},\mathbf{y}^{(j)}\bigr{)}+g\bigl{(}\mathbf{y}^{(j)},\mathbf{y}^{(i)}\bigr{)}\bigr{)}/2$. The Laplacian for the matrix $G$ is defined as $L=I-D^{-\frac{1}{2}}GD^{-\frac{1}{2}}$, where $D$ is a diagonal matrix with elements $D_{ii}=\sum_{j=1}^{M}g_{ij}$. Consequently, the following formula is derived:
$U_{\text{EigV}}=\sum_{i=1}^{M}\max(0,1-\lambda_{i}).$ (18)
@@
Both Number of Semantic Sets and Sum of Eigenvalues of the Graph Laplacian effectively capture overall variation in generated text but cannot produce an individual uncertainty score for each output. To address this, Lin et al. (2024) proposes to use the diagonal Degree Matrix $D$ which represents the total similarity of each answer with all others. The corrected trace of $D$ provides an average pairwise distance between answers, and uncertainty is computed as:
$U_{\mathrm{DegMat}}=1-\mathrm{trace}(D)/M^{2}.$ (19)
C.3 Information-Based Methods with Repeated Sampling
In this section we detail methods that integrate model confidence with consistency.
@@ C.3 Information-Based Methods with Repeated Sampling
We can compute the entropy on the sequence level $\mathbb{E}\bigl{[}-\log P(\mathbf{y}\mid\mathbf{x})\bigr{]}$, where the expectation is taken over the sequences $\mathbf{y}$ randomly generated from the distribution $P(\mathbf{y}\mid\mathbf{x})$. Unfortunately, while for token level, we have an exact way of computing the entropy, for the sequence level, we need to adhere to some approximations. In practice, we can use Monte-Carlo integration, i.e. generate several sequences $\mathbf{y}^{(i)},\,i=1,\dots,M$ via random sampling and compute Monte Carlo Sequence Entropy:
@@
Semantic Entropy (Kuhn et al., 2023) addresses the issue of generated sequences with similar meanings but differing probabilities according to the model, which can heavily influence the resulting entropy value (20). The method clusters generated sequences $\mathbf{y}^{(i)},\,i=1,\dots,M$ into semantically homogeneous groups $\mathcal{C}_{k},~{}k=1,\dots,K$ (where $K\leq M$) using a bi-directional entailment algorithm. Probabilities of sequences are averaged within each cluster. The entropy estimate is then defined as:
@@
SentenceSAR (Duan et al., 2024) enhances the probability of sentences that are more relevant. It uses a sentence relevance measure $g\bigl{(}\mathbf{y}^{(j)},\mathbf{y}^{(k)}\bigr{)}$ to evaluate the relevance of $\mathbf{y}^{(j)}$ with respect to $\mathbf{y}^{(k)}$. SentenceSAR is calculated as:
@@
$\mathrm{R}_{S}(\mathbf{y}^{(j)},\mathbf{x})\!=\sum_{k\neq j}g\bigl{(}\mathbf{y}^{(j)},\mathbf{y}^{(k)}\bigr{)}P\bigl{(}\mathbf{y}^{(k)}\mid\mathbf{x}\bigr{)}.$ (23)
The combination of SentenceSAR and TokenSAR results in a unified method called SAR (Duan et al., 2024). In this approach, the generative probability $P(\mathbf{y}\mid\mathbf{x})$ in the SentenceSAR formula is replaced with the token-shifted probability $P^{\prime}(\mathbf{y}\mid\mathbf{x})=\exp\bigl{\{}-\mathrm{TokenSAR}(\mathbf{y},\mathbf{x})\bigr{\}}$, creating a comprehensive measure that integrates both sentence- and token-level adjustments.
Appendix D Detailed Experimental Results
In this section, we present detailed experimental results, which were used for computing values in Tables 1 and 2.

### v3

@@ C.2 Different Ways of Combining Confidence and Consistency
$U_{\text{FullSampleCoCoA}}(\mathbf{y}_{*}\mid\mathbf{x})=u(\mathbf{y}_{*}\mid\mathbf{x})\cdot U_{\text{DegMat}}(\mathbf{x}).$ (17)
In addition, we present results where we compute $u(\mathbf{y}_{*}\mid\mathbf{x})$ using probability-based scores instead of log-likelihood, see equation (6). We refer to this variant as $U_{\text{ProbCoCoA}}(\mathbf{y}_{*}\mid\mathbf{x})$. While the results are generally similar, the log-likelihood-based formulation offers greater numerical stability. Therefore, we adopt it as the default in the main paper.
@@
XSUM WMT14FrEn WMT19DeEn CoQA TriviaQA MMLU GSM8k
@@
Table 8: Comparison of PRRs of CoCoA-family methods with alternative formulations. Main model response obtained via greedy decoding.
@@
XSUM WMT14FrEn WMT19DeEn CoQA TriviaQA MMLU GSM8k
@@
Table 9: Comparison of PRRs of CoCoA-family methods with alternative formulations. Main model response obtained by selecting the most probable sample.
Method Dataset
XSUM WMT14FrEn WMT19DeEn CoQA TriviaQA MMLU GSM8k
Mistral7b-Base
$\text{AdditiveCoCoA}_{MSP}$ 0.259 0.382 0.488 0.231 0.388 0.512 0.711
$\text{FullSampleCoCoA}_{MSP}$ 0.287 0.447 0.591 0.318 0.570 0.516 0.748
$\text{ProbCoCoA}_{MSP}$ 0.258 0.374 0.473 0.149 0.099 -0.265 0.710
$\text{CoCoA}_{MTE}$ 0.295 0.441 0.589 0.336 0.597 0.514 0.766
$\text{AdditiveCoCoA}_{PPL}$ 0.234 0.430 0.560 0.255 0.517 0.507 0.721
$\text{FullSampleCoCoA}_{PPL}$ 0.251 0.458 0.571 0.305 0.587 0.516 0.745
$\text{ProbCoCoA}_{PPL}$ 0.094 -0.123 -0.143 -0.202 -0.404 -0.423 0.027
$\text{CoCoA}_{PPL}$ 0.258 0.451 0.577 0.322 0.612 0.514 0.768
$\text{AdditiveCoCoA}_{MTE}$ 0.201 0.468 0.585 0.296 0.669 0.501 0.760
$\text{FullSampleCoCoA}_{MTE}$ 0.232 0.426 0.544 0.318 0.673 0.496 0.772
$\text{CoCoA}_{MTE}$ 0.236 0.419 0.553 0.337 0.680 0.494 0.783
Llama8b-Base
$\text{AdditiveCoCoA}_{MSP}$ 0.197 0.345 0.465 0.177 0.287 0.571 0.676
$\text{FullSampleCoCoA}_{MSP}$ 0.214 0.409 0.589 0.289 0.449 0.589 0.724
$\text{ProbCoCoA}_{MSP}$ 0.196 0.337 0.449 0.102 0.053 0.434 0.673
$\text{CoCoA}_{MTE}$ 0.225 0.403 0.590 0.304 0.487 0.586 0.731
$\text{AdditiveCoCoA}_{PPL}$ 0.182 0.406 0.500 0.229 0.401 0.540 0.715
$\text{FullSampleCoCoA}_{PPL}$ 0.186 0.433 0.563 0.304 0.461 0.565 0.746
$\text{ProbCoCoA}_{PPL}$ 0.101 -0.097 -0.152 -0.175 -0.367 0.099 0.053
$\text{CoCoA}_{PPL}$ 0.198 0.428 0.580 0.320 0.496 0.562 0.751
$\text{AdditiveCoCoA}_{MTE}$ 0.160 0.437 0.554 0.302 0.560 0.477 0.748
$\text{FullSampleCoCoA}_{MTE}$ 0.174 0.400 0.543 0.334 0.568 0.475 0.762
$\text{CoCoA}_{MTE}$ 0.189 0.398 0.561 0.345 0.587 0.472 0.765
Falcon7b-Base
$\text{AdditiveCoCoA}_{MSP}$ 0.188 0.356 0.401 0.228 0.488 0.578 0.368
$\text{FullSampleCoCoA}_{MSP}$ 0.237 0.468 0.586 0.325 0.593 0.576 0.453
$\text{ProbCoCoA}_{MSP}$ 0.184 0.339 0.378 0.144 0.350 -0.345 0.364
$\text{CoCoA}_{MTE}$ 0.248 0.461 0.590 0.336 0.617 0.575 0.472
$\text{AdditiveCoCoA}_{PPL}$ 0.206 0.404 0.565 0.271 0.537 0.572 0.625
$\text{FullSampleCoCoA}_{PPL}$ 0.231 0.464 0.606 0.337 0.598 0.577 0.650
$\text{ProbCoCoA}_{PPL}$ 0.014 -0.167 -0.265 -0.187 -0.237 -0.416 -0.206
$\text{CoCoA}_{PPL}$ 0.242 0.462 0.613 0.345 0.624 0.575 0.662
$\text{AdditiveCoCoA}_{MTE}$ 0.191 0.488 0.612 0.353 0.644 0.564 0.617
$\text{FullSampleCoCoA}_{MTE}$ 0.219 0.452 0.591 0.368 0.654 0.556 0.629
$\text{CoCoA}_{MTE}$ 0.237 0.454 0.600 0.373 0.666 0.557 0.650
Table 10: Comparison of PRRs of CoCoA-family methods with alternative formulations. Main model response obtained by MBR decoding.
Appendix D Detailed Description of Uncertainty Quantification Methods
@@
D.1 Information-Based Methods
@@
Maximum Sequence Probability (MSP) is one of the simplest and most direct methods for estimating uncertainty. It measures the probability of the most likely output sequence given a specific input. Thus, uncertainty is quantified by calculating the probability of the sequence with the highest likelihood, under the assumption that the model is most confident in this output. In its simplest form it is given by (6). An equivalent (in terms of ordering of predictions for different inputs) but more numerically stable formulation is logarithmic. It is defined as:
$U_{\text{MSP}}(\mathbf{y}\mid\mathbf{x})=-\log p(\mathbf{y}\mid\mathbf{x}).$ (18)
Perplexity (PPL) is another widely used metric for estimating uncertainty in language models [Fomicheva et al., 2020]. It measures the model’s confidence by evaluating the average likelihood of generating the sequence tokens:
$U_{\mathrm{PPL}}(\mathbf{y}\mid\mathbf{x})=-\frac{1}{L}\log p(\mathbf{y}\mid\mathbf{x}).$ (19)
Mean Token Entropy takes a broader view of uncertainty by considering the token-level predictions across the entire sequence [Fomicheva et al., 2020]. Instead of evaluating the model’s confidence in a single output or individual token predictions, Mean Token Entropy calculates the average entropy of the token probability distributions for each token in the sequence:
$U_{\mathcal{H}_{T}}(\mathbf{y}\mid\mathbf{x})=\frac{1}{L}\sum_{l=1}^{L}\mathcal{H}(y_{l}\mid\mathbf{y}_{<l},\mathbf{x}),$ (20)
@@
The TokenSAR method, introduced in Duan et al. [2024], generalizes length-normalized log probability by computing a weighted average of the negative log probabilities of generated tokens, where weights are based on token relevance to the overall text. Using a similarity function $s(\cdot,\cdot)$ and token relevance function $R_{T}(y_{k},\mathbf{y},\mathbf{x})=1-s(\mathbf{x}\cup\mathbf{y},\mathbf{x}\cup\mathbf{y}\setminus y_{k})$, the uncertainty estimate is calculated as:
$U_{\mathrm{TokenSAR}}(\mathbf{y}\mid\mathbf{x})=-\sum_{l=1}^{L}\tilde{\mathrm{R}}_{T}(y_{l},\mathbf{y},\mathbf{x})\log p(y_{l}\mid\mathbf{y}_{<l},\mathbf{x}),$ (21)
@@
$\tilde{\mathrm{R}}_{T}(y_{k},\mathbf{y}\mid\mathbf{x})=\frac{\mathrm{R}_{T}(y_{k},\mathbf{y},\mathbf{x})}{\sum\nolimits_{l=1}^{L}\mathrm{R}_{T}(y_{l},\mathbf{y},\mathbf{x})}.$ (22)
@@
D.2 Consistency-Based Methods
@@
Consistency-based methods assess the uncertainty of a language model by evaluating the semantic consistency of its predictions across multiple outputs for the same prompt. The core idea is that semantically similar outputs indicate higher confidence, while diverse or conflicting outputs suggest greater uncertainty. Since language models can express the same meaning in different surface forms, these methods construct a semantic similarity matrix $S=(s_{ij})$, where each entry represents the degree of similarity between pairs of responses. By clustering responses into groups with equivalent meanings, these methods provide a semantic measure of the model’s consistency.
@@
The work Lin et al. [2024] offers two similarity measures to evaluate the similarity of sequences. The first is the Jaccard similarity, which treats sequences as sets of words and calculates the proportion of shared words to the total number of unique words in both sequences: $s(\mathbf{y},\mathbf{y}^{\prime})=|\mathbf{y}\cap\mathbf{y}^{\prime}|/|\mathbf{y}\cup\mathbf{y}^{\prime}|$.
@@
Natural Language Inference (NLI) provides another method for computing similarity between sequences. We use the DeBERTa-large NLI model He et al. [2021], following Kuhn et al. [2023]. For each pair of sequences, an NLI model predicts two probabilities: ${p}_{\mathrm{entail}}(\mathbf{y},\mathbf{y}^{\prime})$, indicating entailment, and ${p}_{\mathrm{contra}}(\mathbf{y},\mathbf{y}^{\prime})$, indicating contradiction. Similarity is then defined as either $s_{\mathrm{entail}}(\mathbf{y},\mathbf{y}^{\prime})={p}_{\mathrm{entail}}(\mathbf{y},\mathbf{y}^{\prime})$ or $s_{\mathrm{contra}}(\mathbf{y},\mathbf{y}^{\prime})=1-{p}_{\mathrm{contra}}(\mathbf{y},\mathbf{y}^{\prime})$.
@@
Among the simplest consistency-based approaches are the Number of Semantic Sets and the Sum of Eigenvalues of the Graph Laplacian [Lin et al., 2024]. Number of Semantic Sets estimates how many distinct “meanings” the model produces by clustering its outputs with an NLI model. The number of semantic sets is initially equal to the total number of generated answers, $M$. Two sentences are grouped into the same cluster if the following conditions are satisfied: ${p}_{\text{entail}}\bigl{(}\mathbf{y}^{(i)},\mathbf{y}^{(j)}\bigr{)}>{p}_{\text{contra}}\bigl{(}\mathbf{y}^{(i)},\mathbf{y}^{(j)}\bigr{)}\quad\text{and}\quad{p}_{\text{entail}}\bigl{(}\mathbf{y}^{(j)},\mathbf{y}^{(i)}\bigr{)}>{p}_{\text{contra}}\bigl{(}\mathbf{y}^{(j)},\mathbf{y}^{(i)}\bigr{)}$. This computation is performed for all pairs of answers, and the final number of distinct clusters is denoted by $U_{\text{NumSemSets}}$.
@@
Sum of Eigenvalues of the Graph Laplacian examines global diversity: it constructs a similarity matrix among the sampled outputs and computes a continuous uncertainty score from the eigenvalues of the Laplacian of that similarity graph. The work Lin et al. [2024] proposes computing an averaged similarity matrix as $s_{ij}=\bigl{(}s\bigl{(}\mathbf{y}^{(i)},\mathbf{y}^{(j)}\bigr{)}+s\bigl{(}\mathbf{y}^{(j)},\mathbf{y}^{(i)}\bigr{)}\bigr{)}/2$. The Laplacian for the matrix $S$ is defined as $L=I-D^{-\frac{1}{2}}SD^{-\frac{1}{2}}$, where $D$ is a (diagonal) degree matrix with elements $D_{ii}=\sum_{j=1}^{M}s_{ij}$. Consequently, the following formula is derived:
$U_{\text{EigV}}(\mathbf{x})=\sum_{i=1}^{M}\max(0,1-\lambda_{i}(\mathbf{x})).$ (23)
@@
Both Number of Semantic Sets and Sum of Eigenvalues of the Graph Laplacian effectively capture overall variation in generated text but cannot produce an individual uncertainty score for each output. To address this, the work Lin et al. [2024] proposes to use the diagonal Degree Matrix $D(\mathbf{x})$ which represents the total similarity of each answer with all others. The corrected trace of $D(\mathbf{x})$ provides an average pairwise distance between answers, and uncertainty is computed as:
$U_{\mathrm{DegMat}}(\mathbf{x})=1-\mathrm{trace}(D(\mathbf{x}))/M^{2}.$ (24)
D.3 Information-Based Methods with Repeated Sampling
Averaging uncertainties.
@@ D.3 Information-Based Methods with Repeated Sampling
We can compute the entropy on the sequence level $\mathbb{E}\bigl{[}-\log p(\mathbf{y}\mid\mathbf{x})\bigr{]}$, where the expectation is taken over the sequences $\mathbf{y}$ randomly generated from the distribution $p(\mathbf{y}\mid\mathbf{x})$. Unfortunately, while for token level, we have an exact way of computing the entropy, for the sequence level, we need to adhere to some approximations. In practice, we can use Monte-Carlo integration, i.e. generate several sequences $\mathbf{y}^{(i)},\,i=1,\dots,M$ via random sampling and compute Monte Carlo Sequence Entropy [Kuhn et al., 2023]:
@@
While simple averaging represents a natural way to aggregate uncertainties, it has certain issues related to the nature of LLMs. First of all, in the vast majority of applications, an LLM-based system should produce a single output $\mathbf{y}_{*}$ for an input query. When we consider $U_{\mathcal{H}_{S}}(\mathbf{x})$ or other similar measure, we essentially perform averaging of uncertainties of different sequences, thus somewhat assessing the uncertainty related to the entire generative distribution $p(\mathbf{y}\mid\mathbf{x})$ for the input $\mathbf{x}$, but not for a particular generated sequence $\mathbf{y}_{*}$. This averaged uncertainty might not be adequate for this particular sequence and, remarkably, often performs worse than the uncertainty $u_{*}=U(\mathbf{y}_{*}\mid\mathbf{x})$, which is related solely to the output $\mathbf{y}_{*}$. Moreover, although intuitive, this naïve aggregation method assumes that all outputs contribute equally to the final uncertainty estimate, regardless of their semantic relationships. This can lead to inconsistencies when semantically equivalent outputs have varying uncertainty scores or when outputs with low similarity are treated as equally important.
Semantically weighted averaging. Semantic Entropy Kuhn et al. [2023] addresses the issue of generated sequences with similar meanings but differing probabilities according to the model, which can heavily influence the resulting entropy value (25). The method clusters generated sequences $\mathbf{y}^{(i)},\,i=1,\dots,M$ into semantically homogeneous groups $\mathcal{C}_{k},~{}k=1,\dots,K$ (where $K\leq M$) using a bi-directional entailment algorithm. Probabilities of sequences are averaged within each cluster. The entropy estimate is then defined as:
@@
SentenceSAR Duan et al. [2024] enhances the probability of sentences that are more relevant. It uses a sentence relevance measure $s\bigl{(}\mathbf{y}^{(j)},\mathbf{y}^{(k)}\bigr{)}$ to evaluate the relevance of $\mathbf{y}^{(j)}$ with respect to $\mathbf{y}^{(k)}$. SentenceSAR is calculated as:
@@
$\mathrm{R}_{S}(\mathbf{y}^{(j)},\mathbf{x})\!=\sum_{k\neq j}s\bigl{(}\mathbf{y}^{(j)},\mathbf{y}^{(k)}\bigr{)}p\bigl{(}\mathbf{y}^{(k)}\mid\mathbf{x}\bigr{)}.$ (28)
The combination of SentenceSAR and TokenSAR results in a unified method called SAR Duan et al. [2024]. In this approach, the generative probability $p(\mathbf{y}\mid\mathbf{x})$ in the SentenceSAR formula is replaced with the token-shifted probability $p^{\prime}(\mathbf{y}\mid\mathbf{x})=\exp\bigl{\{}-\mathrm{TokenSAR}(\mathbf{y},\mathbf{x})\bigr{\}}$, creating a comprehensive measure that integrates both sentence- and token-level adjustments.
The aggregation approaches like Semantic Entropy Kuhn et al. [2023] or SAR Duan et al. [2024] can be unified into a semantically-aware Generalized Monte Carlo uncertainty estimate, defined as
$U_{\text{GMCU}}(\mathbf{x})=\frac{1}{M}\sum_{i=1}^{M}h\Biggl{(}\sum_{j=1}^{M}s\bigl{(}\mathbf{y}^{(i)},\mathbf{y}^{(j)}\bigr{)}p\bigl{(}\mathbf{y}^{(j)}\mid\mathbf{x}\bigr{)}\Biggr{)}.$ (29)
Here, the inner summation aggregates sequence probabilities $p\bigl{(}\mathbf{y}^{(j)}\mid\mathbf{x}\bigr{)}$ weighted by their semantic similarity to the $i$-th output, and the outer summation averages these contributions across all samples. The function $h(\cdot)$ provides an additional layer of flexibility, transforming the reweighted uncertainty scores, making the method a generalized framework for uncertainty quantification.
Unfortunately, methods that fall under GMCU, while offering benefits, also inherit the aforementioned issues from both categories of methods. In particular, the outer summation in (29), similarly to the case of simple Monte Carlo averaging, often fails to outperform the uncertainty $U_{*}(\mathbf{x})=U(\mathbf{y}_{*}\mid\mathbf{x})$ of a single generated sequence $\mathbf{y}_{*}$.
Appendix E Detailed Experimental Results
Tables 3 and 1 presented PRR scores averaged over datasets corresponding to each task. Here we present expanded results for each dataset, see Tables 11, 12 and 13.